# Towards Generalizable Tumor Synthesis

### v1

@@ 1 Introduction
Tumor synthesis enables the creation of artificial tumor examples in medical images, facilitating the training of AI models for tumor detection and segmentation [jordon2018pate, yoon2019time, chen2021synthetic]. Synthetic tumors are particularly valuable when there is a scarcity of per-voxel annotated real tumors for effective AI training. Typically, to train AI models for tumor detection in multiple ($N$) organs, annotated real tumor examples from each of these organs are necessary, and ideally, in substantial numbers [li2024well, kang2023label, liu2023clip, chou2024acquiring]. Furthermore, AI models often fail to generalize across images from different hospitals, which may vary due to variations in imaging protocols, patient demographics, and scanner manufacturers [orbes2019multi, zhou2022interpreting, zhou2021towards]. The challenge amplifies with the need for extensive manual annotations, a task that could demand up to 25 human years for annotating just one tumor type [xia2022felix, chen2023cancerunit, abi2023automatic]. The task of collecting and annotating a comprehensive dataset encompassing tumors from multiple organs ($N$) and images from numerous hospitals ($M$) is daunting, considering both annotation cost and complexity ($N$$\times$$M$). We hypothesize that tumor synthesis could solve this challenge by creating various tumor types across non-tumor images from multiple hospitals, even when only one tumor type is available, thereby simplifying the complexity from $N$$\times$$M$ to $1$$\times$$M$.
@@ 4.1 Visual Turing Test
We conduct the Visual Turing Test on 240 CT volumes for three organs, respectively, where 120 volumes are with real tumors, and the remaining 120 volumes are with synthesized tumors by our method. Four radiologists, with varying levels of experience ranging from junior to senior and professional, are involved in this test. The total Visual Turing Test took 144 hours (2,880 CTs). Following Hu et al. [hu2023label], each sample is inspected in a 3D view to be classified as either real or synthetic, allowing for the observation of a continuous slice sequence. The testing results are shown in Table 1. Radiologists R1 and R2 can both identify real tumors with a high sensitivity score (above 90%). This indicates their familiarity with the characteristics of real tumors. However, R1’s near-zero specificity scores for liver and pancreas indicate that the synthetic data strongly resembles real tumors, leading to most synthetic tumors being misidentified as real ones. This results in R1’s accuracy scores hovering around 50%. As for R2, who has more experience, the specificity scores are higher than that of R1, approximating 50%. This suggests that nearly 50% of synthetic samples are still incorrectly identified as real samples. These results confirm the efficacy of DiffTumor in generating visually realistic tumors.
R1
R2
R3
R4
liver
sensitivity (%)
100
99.2
specificity (%)
2.9
53.3
accuracy (%)
45.0
76.3
pancreas
sensitivity (%)
97.1
100
specificity (%)
0.0
44.2
accuracy (%)
56.7
72.1
kidney
sensitivity (%)
95.8
98.3
99.2
97.5
specificity (%)
36.7
55.0
40.8
51.7
accuracy (%)
66.3
76.7
70.0
74.6
@@ 4.2 Generalizable to Different Organs
liver
pancreas
kidneys
liver
real tumors
75.6
0
2.4
Hu et al. [hu2023label]
77.8
56.3
52.4
DiffTumor
82.2 56.3 76.2
pancreas
real tumors
0.7
64.3
0
Hu et al. [hu2023label]
74.1
67.0
52.4
DiffTumor
75.3 71.4 71.4
kidney
real tumors
0.1
0
50.0
Hu et al. [hu2023label]
74.1
56.3
66.7
DiffTumor
68.8
61.6 78.6
@@
backbone
method
liver
pancreas
kidneys
U-Net
real tumors
62.3$\pm$28.3
56.0$\pm$24.8
75.1$\pm$27.2
DiffTumor
70.9$\pm$21.1 64.8$\pm$24.5 84.2$\pm$9.5
nnU-Net
real tumors
64.3$\pm$26.5
59.9$\pm$23.8
73.8$\pm$20.9
DiffTumor
73.6$\pm$18.1 63.6$\pm$27.7 84.5$\pm$11.5
SwinUNETR
real tumors
65.1$\pm$23.5
52.2$\pm$31.2
80.6$\pm$19.6
DiffTumor
71.4$\pm$19.1 62.2$\pm$26.1 85.1$\pm$8.7
@@ 5 Related Work
Generative models such as Energy-Based Models [lecun2006tutorial, zhao2016energy, du2019implicit], Variational Autoencoders (VAE) [kingma2013auto, kingma2014semi, kingma2019introduction], Generative Adversarial Networks (GAN) [goodfellow2014generative, goodfellow2016deep, goodfellow2020generative, creswell2018generative, jordon2018pate, yoon2019time], and normalizing flows [papamakarios2021normalizing, kobyzev2020normalizing, yu2021fastflow] have shown significant potential in creating realistic images. Among these, Diffusion Models [sohl2015deep, ho2020denoising, song2020score] and their variants [vahdat2021score, kim2021maximum, rombach2022high] have recently emerged as particularly advanced in image generation. In the medical field, generative models have been effectively utilized for tasks like image-to-image translation [lyu2022conversion, meng2022novel, ozbey2023unsupervised], reconstruction [song2021solving, xie2022measurement], segmentation [fernandez2022can, kim2022diffusion, wolleb2022diffusion], image denoising [gong2023pet], and anomaly detection [wyatt2022anoddpm, siddiquee2019learning, xiang2023squid]. In this work, we focus on generating tumors in abdominal organs based on the textures of the surrounding organs, which significantly reduces the annotated data required for training. Refer to Appendix F for a more comprehensive comparison with our DiffTumor.
@@
Tumor synthesis that is widely effective for a variety of organs is an attractive topic. Successful works related to tumor synthesis based on various medical modalities include colon polyp synthesis in colonoscopy videos [shin2018abnormal], tumor cell synthesis in fluorescence microscopy images [horvath2022metgan], synthesized brain tumors in MRI [billot2023synthseg] for pre-training [zhang2023self], lung nodule synthesis in CT images [han2019synthesizing, yang2019class, jin2021free], and lesion in dermatoscopic images [du2023boosting]. Additionally, many works on synthesizing non-cancerous lesions have been explored, such as COVID-19 lesion synthesis in chest CT [lyu2022pseudo, yao2021label], and diabetic lesion synthesis in retinal images [wang2022anomaly]. Recent studies have improved the realism of synthetic tumors in the liver [zhang2023unsupervised, lyu2022learning, hu2022synthetic, hu2023synthetic] and pancreas [wei2022pancreatic, li2023early]. AI trained on these synthetic tumors perform similarly well as those trained with real tumors. However, these methods need to be redesigned for tumors in other organs, which severely limits the generalization capabilities. In this work, we learn the tumor distribution based on generative models, i.e., Diffusion Models, to realize generalizable tumor synthesis.
@@ Appendix C Generalizable to Multiple Organs
liver
pancreas
kidneys
liver
real tumors
77.4
1.8
2.4
Hu et al. [hu2023label]
78.0
56.3
61.9
DiffTumor
80.9 60.7 71.4
pancreas
real tumors
1.5
67.0
2.4
Hu et al. [hu2023label]
76.2
68.8
61.9
DiffTumor
72.8
75.0 81.0
kidney
real tumors
0.9
0.9
59.5
Hu et al. [hu2023label]
76.2
56.3
69.0
DiffTumor
77.4 63.4 76.2
@@
liver
pancreas
kidneys
liver
real tumors
76.2
2.7
0
Hu et al. [hu2023label]
79.2
63.4
71.4
DiffTumor
83.1 69.6 73.8
pancreas
real tumors
1.6
70.5
4.8
Hu et al. [hu2023label]
66.4
73.2
71.4
DiffTumor
76.2 79.5 90.5
kidney
real tumors
0.6
0
69.0
Hu et al. [hu2023label]
66.4
63.4
76.2
DiffTumor
76.2
61.6
81.0
@@
U-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
62.3
72.8
63.8
54.5
59.0
62.5
NSD (%)
63.4
74.6
63.3
55.5
61.6
63.7
DiffTumor
DSC (%)
70.9
74.0
67.9
59.1
60.6
66.5
NSD (%)
71.2
73.9
70.4
61.0
63.3
68.0
nnU-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
64.3
70.2
64.3
56.3
59.6
62.9
NSD (%)
65.7
72.7
63.1
59.3
62.6
64.7
DiffTumor
DSC (%)
73.6
73.9
67.6
64.9
63.8
68.8
NSD (%)
75.3
73.9
67.9
69.0
66.5
70.5
Swin UNETR
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
65.1
69.4
57.4
59.0
58.2
61.8
NSD (%)
65.9
71.7
53.1
62.1
61.9
62.8
DiffTumor
DSC (%)
71.4
71.7
71.6
62.2
62.4
67.9
NSD (%)
73.5
72.4
74.5
66.5
66.0
70.6
@@
U-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
56.0
51.9
45.5
59.4
43.2
51.2
NSD (%)
51.0
49.9
43.6
57.7
40.2
48.5
DiffTumor
DSC (%)
64.8
58.0
57.7
67.9
51.8
60.0
NSD (%)
60.5
55.3
58.3
67.5
47.9
57.9
nnU-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
59.9
50.0
44.8
63.5
50.4
53.7
NSD (%)
55.7
47.0
47.0
62.3
48.0
52.0
DiffTumor
DSC (%)
63.6
60.5
62.5
67.8
55.3
61.9
NSD (%)
61.1
59.1
63.4
67.7
54.9
61.2
Swin UNETR
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
52.2
49.5
50.9
60.0
52.1
52.9
NSD (%)
49.1
49.7
50.9
58.9
47.4
51.2
DiffTumor
DSC (%)
62.2
60.2
59.0
69.7
53.8
61.0
NSD (%)
58.7
58.4
62.8
67.2
51.2
59.7
@@
U-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
75.1
68.0
69.0
78.1
70.6
72.0
NSD (%)
68.4
59.0
57.7
68.3
62.4
63.2
DiffTumor
DSC (%)
84.2
76.7
79.4
80.6
74.1
79.0
NSD (%)
76.6
64.5
70.7
71.7
65.8
69.9
nnU-Net
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
73.8
76.8
80.0
80.5
73.4
76.9
NSD (%)
62.7
70.2
71.2
70.8
67.5
68.5
DiffTumor
DSC (%)
84.5
83.4
81.6
83.9
77.3
82.1
NSD (%)
78.3
74.4
74.1
76.9
72.3
75.2
Swin UNETR
metrics
fold0
fold1
fold2
fold3
fold4
average
real tumors
DSC (%)
80.6
64.9
79.1
76.0
72.2
74.6
NSD (%)
74.2
55.5
68.2
67.4
64.8
66.0
DiffTumor
DSC (%)
85.1
77.2
81.2
85.7
79.9
81.8
NSD (%)
79.2
70.8
74.1
78.0
74.1
75.2
@@ F.1 Comparison with Related Works
DSC (%)
NSD (%)
DSC (%)
NSD (%)
DSC (%)
NSD (%)
real tumors
62.3
63.4
56.0
51.0
75.1
68.4
Hu et al. [hu2023label]
69.7
70.9
55.9
49.9
80.8
71.0
DiffTumor
70.9 71.2 64.8 60.5 84.2 76.6

### v2

@@ 1 Introduction
Tumor synthesis enables the creation of artificial tumor examples in medical images [jordon2018pate, yoon2019time, chen2021synthetic], it is particularly valuable when there is a dearth or complete absence of per-voxel annotated real tumors (e.g., early-stage tumors) for effective AI training. Typically, to train AI models for tumor detection in multiple ($N$) organs, annotated real tumor examples from each of these organs are necessary, and ideally, in substantial numbers [li2024well, kang2023label, liu2023clip, chou2024acquiring, zhu2022assembling]. Furthermore, AI models often fail to generalize across images from different hospitals, which may vary due to variations in imaging protocols, patient demographics, and scanner manufacturers [orbes2019multi, zhou2022interpreting, zhou2021towards]. The challenge amplifies with the need for extensive manual annotations, a task that could demand up to 25 human years for annotating just one tumor type [xia2022felix, chen2023cancerunit, abi2023automatic]. The task of collecting and annotating a comprehensive dataset encompassing tumors from multiple organs ($N$) and images from numerous hospitals ($M$) is daunting, considering both annotation cost and complexity ($N$$\times$$M$). We hypothesize that tumor synthesis could solve this challenge by creating various tumor types across non-tumor images from multiple hospitals, even when only one tumor type is available, thereby simplifying the complexity from $N$$\times$$M$ to $1$$\times$$M$.
@@ 4.1 Visual Turing Test
We conduct the Visual Turing Test on 240 CT volumes for three organs, respectively, where 120 volumes are with real tumors, and the remaining 120 volumes are with synthesized tumors by our method. Four radiologists, with varying levels of experience ranging from junior to senior and professional, are involved in this test. The total Visual Turing Test took 144 hours (2,880 CTs). Following Hu et al. [hu2023label], each sample is inspected in a 3D view to be classified as either real or synthetic, allowing for the observation of a continuous slice sequence. The testing results are shown in Table 1. All radiologists are able to identify real tumors with a high sensitivity score (above 90%). This indicates their familiarity with the characteristics of real tumors. However, the low specificity scores (below 40%) on the three types of tumors for radiologists R1 and R3 suggest that the synthetic data strongly resembles real tumors, leading to most synthetic tumors being misidentified as real ones. As for R2 and R4, who have more experience, the specificity scores are higher than those of R1 and R3, approximating 50%. This indicates that nearly 50% of synthetic samples are still incorrectly identified as real tumors. These results confirm the efficacy of DiffTumor in generating visually realistic tumors.
R1 R2 R3 R4
liver sensitivity (%) 98.3 99.2 100 100
specificity (%) 31.7 53.3 39.2 45.8
accuracy (%) 65.0 76.3 69.6 72.9
pancreas sensitivity (%) 96.7 100 100 98.3
specificity (%) 22.5 44.2 34.2 38.8
accuracy (%) 59.6 72.1 67.1 68.3
kidney sensitivity (%) 95.8 98.3 99.2 97.5
specificity (%) 36.7 55.0 40.8 51.7
accuracy (%) 66.3 76.7 70.0 74.6
@@ 4.2 Generalizable to Different Organs
liver real tumors 75.6 0 2.4
Hu et al. [hu2023label] 77.8 56.3 52.4
DiffTumor 82.2 56.3 76.2
pancreas real tumors 0.7 64.3 0
Hu et al. [hu2023label] 74.1 67.0 52.4
DiffTumor 75.3 71.4 71.4
kidney real tumors 0.1 0 50.0
Hu et al. [hu2023label] 74.1 56.3 66.7
DiffTumor 68.8 61.6 78.6
@@
backbone method liver pancreas kidneys
U-Net real tumors 62.3$\pm$28.3 56.0$\pm$24.8 75.1$\pm$27.2
DiffTumor 70.9$\pm$21.1 64.8$\pm$24.5 84.2$\pm$9.5
nnU-Net real tumors 64.3$\pm$26.5 59.9$\pm$23.8 73.8$\pm$20.9
DiffTumor 73.6$\pm$18.1 63.6$\pm$27.7 84.5$\pm$11.5
SwinUNETR real tumors 65.1$\pm$23.5 52.2$\pm$31.2 80.6$\pm$19.6
DiffTumor 71.4$\pm$19.1 62.2$\pm$26.1 85.1$\pm$8.7
@@ 5 Related Work
Generative models such as Energy-Based Models [lecun2006tutorial, zhao2016energy, du2019implicit], Variational Autoencoders (VAE) [kingma2013auto, kingma2014semi, kingma2019introduction], Generative Adversarial Networks (GAN) [goodfellow2014generative, goodfellow2016deep, goodfellow2020generative, creswell2018generative, jordon2018pate, yoon2019time, chen2022mask], and normalizing flows [papamakarios2021normalizing, kobyzev2020normalizing, yu2021fastflow] have shown significant potential in creating realistic images. Among these, Diffusion Models [sohl2015deep, ho2020denoising, song2020score] and their variants [vahdat2021score, kim2021maximum, rombach2022high] have recently emerged as particularly advanced in image generation. In the medical field, generative models have been effectively utilized for tasks like image-to-image translation [lyu2022conversion, meng2022novel, ozbey2023unsupervised], reconstruction [song2021solving, xie2022measurement], segmentation [fernandez2022can, kim2022diffusion, wolleb2022diffusion], image denoising [gong2023pet], and anomaly detection [wyatt2022anoddpm, siddiquee2019learning, xiang2023squid]. In this work, we focus on generating tumors in abdominal organs based on the textures of the surrounding organs, which significantly reduces the annotated data required for training. Refer to Appendix F for a more comprehensive comparison with our DiffTumor.
@@
Tumor synthesis that is widely effective for a variety of organs is an attractive topic. Successful works related to tumor synthesis based on various medical modalities include colon polyp synthesis in colonoscopy videos [shin2018abnormal], tumor cell synthesis in fluorescence microscopy images [horvath2022metgan], synthesized brain tumors [billot2023synthseg, zhang2023self] and myocardial pathology [zhang2024lefusion] in MRI, lung nodule synthesis in CT images [han2019synthesizing, yang2019class, jin2021free], and lesion in dermatoscopic images [du2023boosting]. Additionally, there are many works on synthesizing non-cancerous lesions such as COVID-19 lesion synthesis in chest CT [lyu2022pseudo, yao2021label], and diabetic lesion synthesis in retinal images [wang2022anomaly]. Recent studies have improved the realism of synthetic tumors in the liver [zhang2023unsupervised, lyu2022learning, hu2022synthetic, hu2023synthetic] and pancreas [wei2022pancreatic, li2023early]. AI trained on these synthetic tumors perform similarly well as those trained with real tumors. However, these methods need to be redesigned for tumors in other organs, which severely limits the generalization capabilities. In this work, we learn the tumor distribution based on generative models, i.e., Diffusion Models, to realize generalizable tumor synthesis.
@@ Appendix C Generalizable to Multiple Organs
liver real tumors 77.4 1.8 2.4
Hu et al. [hu2023label] 78.0 56.3 61.9
DiffTumor 80.9 60.7 71.4
pancreas real tumors 1.5 67.0 2.4
Hu et al. [hu2023label] 76.2 68.8 61.9
DiffTumor 72.8 75.0 81.0
kidney real tumors 0.9 0.9 59.5
Hu et al. [hu2023label] 76.2 56.3 69.0
DiffTumor 77.4 63.4 76.2
@@
liver real tumors 76.2 2.7 0
Hu et al. [hu2023label] 79.2 63.4 71.4
DiffTumor 83.1 69.6 73.8
pancreas real tumors 1.6 70.5 4.8
Hu et al. [hu2023label] 66.4 73.2 71.4
DiffTumor 76.2 79.5 90.5
kidney real tumors 0.6 0 69.0
Hu et al. [hu2023label] 66.4 63.4 76.2
DiffTumor 76.2 61.6 81.0
@@
U-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 62.3 72.8 63.8 54.5 59.0 62.5
NSD (%) 63.4 74.6 63.3 55.5 61.6 63.7
DiffTumor DSC (%) 70.9 74.0 67.9 59.1 60.6 66.5
NSD (%) 71.2 73.9 70.4 61.0 63.3 68.0
nnU-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 64.3 70.2 64.3 56.3 59.6 62.9
NSD (%) 65.7 72.7 63.1 59.3 62.6 64.7
DiffTumor DSC (%) 73.6 73.9 67.6 64.9 63.8 68.8
NSD (%) 75.3 73.9 67.9 69.0 66.5 70.5
Swin UNETR metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 65.1 69.4 57.4 59.0 58.2 61.8
NSD (%) 65.9 71.7 53.1 62.1 61.9 62.8
DiffTumor DSC (%) 71.4 71.7 71.6 62.2 62.4 67.9
NSD (%) 73.5 72.4 74.5 66.5 66.0 70.6
@@
U-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 56.0 51.9 45.5 59.4 43.2 51.2
NSD (%) 51.0 49.9 43.6 57.7 40.2 48.5
DiffTumor DSC (%) 64.8 58.0 57.7 67.9 51.8 60.0
NSD (%) 60.5 55.3 58.3 67.5 47.9 57.9
nnU-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 59.9 50.0 44.8 63.5 50.4 53.7
NSD (%) 55.7 47.0 47.0 62.3 48.0 52.0
DiffTumor DSC (%) 63.6 60.5 62.5 67.8 55.3 61.9
NSD (%) 61.1 59.1 63.4 67.7 54.9 61.2
Swin UNETR metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 52.2 49.5 50.9 60.0 52.1 52.9
NSD (%) 49.1 49.7 50.9 58.9 47.4 51.2
DiffTumor DSC (%) 62.2 60.2 59.0 69.7 53.8 61.0
NSD (%) 58.7 58.4 62.8 67.2 51.2 59.7
@@
U-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 75.1 68.0 69.0 78.1 70.6 72.0
NSD (%) 68.4 59.0 57.7 68.3 62.4 63.2
DiffTumor DSC (%) 84.2 76.7 79.4 80.6 74.1 79.0
NSD (%) 76.6 64.5 70.7 71.7 65.8 69.9
nnU-Net metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 73.8 76.8 80.0 80.5 73.4 76.9
NSD (%) 62.7 70.2 71.2 70.8 67.5 68.5
DiffTumor DSC (%) 84.5 83.4 81.6 83.9 77.3 82.1
NSD (%) 78.3 74.4 74.1 76.9 72.3 75.2
Swin UNETR metrics fold0 fold1 fold2 fold3 fold4 average
real tumors DSC (%) 80.6 64.9 79.1 76.0 72.2 74.6
NSD (%) 74.2 55.5 68.2 67.4 64.8 66.0
DiffTumor DSC (%) 85.1 77.2 81.2 85.7 79.9 81.8
NSD (%) 79.2 70.8 74.1 78.0 74.1 75.2
@@ F.1 Comparison with Related Works
DSC (%) NSD (%) DSC (%) NSD (%) DSC (%) NSD (%)
real tumors 62.3 63.4 56.0 51.0 75.1 68.4
Hu et al. [hu2023label] 69.7 70.9 55.9 49.9 80.8 71.0
DiffTumor 70.9 71.2 64.8 60.5 84.2 76.6